Blades Inc., U.S. firm has agreed to sell its product to Regency, EU firm and payment is Euro 2 million. Blades Inc has agreement that the amount will be paid in 30 days. Blades Inc wants to avoid the FX risk by using the US and EU money market. Following information is available to Blades Inc.

Current Spot Rate = $1.15/Euro

EU Borrowing Rate = 6% annual rate

US Investing Rate = 5% annual rate

Explain how blades inc can conduct money market hedge?
Answer #1

Following steps will be undertaken under money market hedge:

Receivables after 30 days = Euro 2,000,000

Borrow such amount in Euro that payable after 30 days equal receivables = 2,000,000/(1+6%*30/360)

= Euro 1,990,049.75

Convert into USD today 1,990,049.75*1.15 = $2,288,557.21

Invest for 30 days and get 2,288,557.21(1+5%*30/360) = $2,298,092.87

Amount in Euro received after 30 days will be used to repay loan in Euro